If the complement of an angle is 6° less than one-third of its supplement, find the measure of the angle.
Answer by nerdybill(7384)

If the complement of an angle is 6° less than one-third of its supplement, find the measure of the angle.
.
Let x = measure of angle
then
90-x = complement
180-x = supplement
.
90-x = (1/3)(180-x)-6
270-3x = (180-x)-18
270-3x = 162-x
270 = 162+2x
270 = 162+2x
108 = 2x
54 degrees = x